OpenStudy (anonymous):

PLS help y=-x^2+2x graph

OpenStudy (anonymous):

the coefficient of the parabola is -1 so it will be opening downwards... it's x-intercepts are x=0 and x=-2. which means the vertex is somewhere where x=-1. can you figure out the y-coordinate of the vertex?

OpenStudy (anonymous):

how did you get -2

OpenStudy (anonymous):

factor the equation: y = -x^2 + 2 = -x*(x - 2) oops. i mean x=2.. sorry.

OpenStudy (anonymous):

still having trouble sorry

OpenStudy (anonymous):

uh hello?

OpenStudy (anonymous):

sorrry still here

OpenStudy (anonymous):

do you understand the factoring?

OpenStudy (anonymous):

so the vertex lies between these two x-intercepts at x=1...

OpenStudy (anonymous):

sorry for being late but i am confused now because your way is different then what the book says

OpenStudy (anonymous):

ok... i tell you what... i'll explain it the way I do it and you can see that we arrive at the same answer...

OpenStudy (anonymous):

ok then

OpenStudy (anonymous):

there is more than one way to do this problem....

OpenStudy (anonymous):

wait doesn't x=-1?

OpenStudy (anonymous):

forget that... that was before you found my mistake. the vertex now should be in the middle of x=0 and x=2. so x=1 is the x-coordinate of the vertex.

OpenStudy (anonymous):

are you clear up to here?

OpenStudy (anonymous):

uh...still confused alittle....

OpenStudy (anonymous):

y = -x^2 + 2x is your equation. i factored the expression so I could find where the graph crosses the x-axis: y = -x(x-2) finding where the graph crosses the x-axis means you replace y with 0: 0 = -x(x-2) obviously, x=0, x=2.

OpenStudy (anonymous):

The way that my book does you find the axis symmetry with x=-b/2a

OpenStudy (anonymous):

i think....

OpenStudy (anonymous):

so you want to do it the way your book does it then?

OpenStudy (anonymous):

yah if it is ok with you

OpenStudy (anonymous):

ok, identify these for me: a = b = c =

OpenStudy (anonymous):

forget c.

OpenStudy (anonymous):

a=-1 b=2 c=0

OpenStudy (anonymous):

so x=-b/(2a) = -2/(2*(-1)) =

OpenStudy (anonymous):

OH!!! i know what i did wrong :P

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

but could you still continue?

OpenStudy (anonymous):

yes... what did we just find? or what is the significance of x=1?

OpenStudy (anonymous):

Huh??

OpenStudy (anonymous):

you just calculated x = -b/(2a)... what's that for? don't say because the book does it..

OpenStudy (anonymous):

is the vertex (1,1)?

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

but how do you find the other four points?

OpenStudy (anonymous):

what other four? your problem just said to graph it...

OpenStudy (anonymous):

oh. i see. you want verification points...

OpenStudy (anonymous):

It says use the vertex and at least 4 other points

OpenStudy (anonymous):

fill out this table for me: